OpenStudy (anonymous):

Find the exact value of the trigonometric function. cot(-5π/4) A. -1 B. -√2 C. 1 D. -√2/2

OpenStudy (jdoe0001):

are you doing trig I assume?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

IS IT A

OpenStudy (anonymous):

\[\tan (-5\pi/4)=-\tan(5\pi/4)=-\tan(\pi/4)=-1 >>>> \cot(-5\pi/4)=1/\tan(-5\pi/4)=1/-1=-1\]

OpenStudy (jdoe0001):

ok, then check your Unit Circle :), for \(\cfrac{5\pi}{4}\) counterclockwise, that is, going BELOW 0, rather than ABOVE 0

OpenStudy (anonymous):

so it's A

OpenStudy (anonymous):

ok thanks guys @ikarus1991 @jdoe0001

OpenStudy (jdoe0001):

yw

OpenStudy (jdoe0001):

actually, it was supposed to be clockwise, but anyhow :|